whitemonsterbunny17 (whitemonsterbunny17):

Ugh, more Geometry. .-. Help with confusing question, please? Thanks!! (ノ^ヮ^)ノ*:・゚✧*:・゚✧

whitemonsterbunny17 (whitemonsterbunny17):

The sum of the measures of the interior angles of a polygon is five times the sum of the measures of its exterior angles, one angle at each vertex. How many sides does the polygon have?

OpenStudy (sleepyjess):

we will use s for interior and o for exterior since \(\it i\) and \(\it e\) are something completely different

OpenStudy (calculusfunctions):

The sum of the interior angles of any polygon is \[180(n -2)\]where n is the # of sides. The sum sum of the exterior angles of any polygon is 360 degrees.

OpenStudy (sleepyjess):

exterior angles must equal 360 so we know that o=360 s=5(360)

OpenStudy (sleepyjess):

s=1800

OpenStudy (sleepyjess):

that is the sum of the interior angles so now we must find the number of sides

OpenStudy (calculusfunctions):

@sleepyjess teach, don't just give the answers. You're really not helping if you're just giving the answers.

whitemonsterbunny17 (whitemonsterbunny17):

Ehh, I don't see where any answers were given @calculusfunctions But okay.. the formula is 180(n-2) Then what? (Meaning what is the next step?)

whitemonsterbunny17 (whitemonsterbunny17):

I know we have to do some form of algebra next, but how do I set it up?

OpenStudy (sleepyjess):

you know that the sum is 1800 so you would do 1800=180(n-2) then solve for n do you know how to start that?

whitemonsterbunny17 (whitemonsterbunny17):

You would divide by 180, right?

OpenStudy (sleepyjess):

Yes, because you are multiplying on the right side, you would divide both sides to get rid of 180.

OpenStudy (sleepyjess):

so what is 1800/180?

whitemonsterbunny17 (whitemonsterbunny17):

I got 10

OpenStudy (sleepyjess):

correct so now what is your equation?

whitemonsterbunny17 (whitemonsterbunny17):

10(n-2), right?

OpenStudy (sleepyjess):

almost you would have 10=n-2 remember you got rid of 180 on the right side

whitemonsterbunny17 (whitemonsterbunny17):

Ohh okay. So now we get rid of the n?

OpenStudy (sleepyjess):

nope now we get rid of the -2 how do you think you would do that?

whitemonsterbunny17 (whitemonsterbunny17):

Add to both sides?

OpenStudy (sleepyjess):

yes so now we have 10+2=n-2+2 which is equal to 12=n

OpenStudy (sleepyjess):

to confirm that you can substitute 12 for n and see if it equals 1800

OpenStudy (sleepyjess):

180(12-2) 180(10) 1800

whitemonsterbunny17 (whitemonsterbunny17):

Ohhhh, okay, that makes sense now!! \(\Large\frak\color{blueviolet}{Thanks!}\)

OpenStudy (sleepyjess):

\(\Large\frak\color{blueviolet}{No~problem!}\)

OpenStudy (sleepyjess):

so our shape has 12 sides \(\Huge\ddot\smile\ddot\smile\ddot\smile\ddot\smile\ddot\smile\ddot\smile\ddot\smile\ddot\smile\ddot\smile\ddot\smile\\\Huge\ddot\smile\ddot\smile\ddot\smile\ddot\smile\)

whitemonsterbunny17 (whitemonsterbunny17):

\(\Large\frak\color{magenta}{YAYYYY!!!!}\) XD

OpenStudy (sleepyjess):

\(\Huge\cal\color{purple}{Don't~worry,\\you~and~I~both~feel~the\\same~way,~you~about\\geometry,~me~about\\pre-calc.}\)